# Building Confidence in Deep Generative Protein Design

### v1

@@ Abstract
Deep generative models show promise for de novo protein design, but their effectiveness within specific protein families remains underexplored. In this study, we evaluate two 3D rigid-body generative methods, score matching and flow matching, to generate monomeric protein backbones in $\mathrm{SE}(3)$ space. Our goal is to provide new insights and build confidence in the broader applicability of deep generative models for protein design. The optimal amino acid sequences were predicted from the generated backbones, followed by side-chain homology modeling. Results demonstrated high structural integrity, with conserved key residues aligning with known proteins. Structural phylogenetic analysis shows evolutionary links between the generated samples and their protein family members. Further molecular dynamics simulations and protein-ligand docking confirm the dynamic stability and functional potential of these samples, with ligand binding inducing conformational changes consistent with those in wild-type proteins.
Machine Learning
@@ 1 Introduction
Protein functions and specificity are determined by their complex structures. Over the past 60 years, we have progressed from viewing protein design as unattainable to achieving complete artificial design and synthesis (Korendovych & DeGrado, 2020; Huang et al., 2016), with expanding applications across industries (Arunachalam et al., 2021; Kingwell, 2024; Barclay & Acharya, 2023). However, precise design remains challenging due to the nonlinear complexity of protein folding and the sensitivity of function to slight changes, still demanding significant resources.
With the rapid growth of protein structure databases (Jamasb et al., 2024; Berman, 2000), various in silico de novo protein design methods, particularly deep generative models (Watson et al., 2023; Ingraham et al., 2023; Wu et al., 2022), have emerged. Approaches such as diffusion-based score matching (SM) (Yim et al., 2023), and flow matching (FM) (Bose et al., 2023), show competitive performance for generating monomeric protein backbones in $\mathrm{SE}(3)$ [^1] Special Euclidean group. See more in Section A.1. space.
@@
While many methods show novelty, diversity, and designability in generated structures, few studies assess their functionality or evolutionary relevance, leaving the observed novelty and diversity potentially as mere hallucinations (Ji et al., 2023). Unclear biological functionality, unknown stability, and the lack of validation connecting these structures to known functions limit the broader application and advancement of generative methods.
This study evaluates the effectiveness of SM and FM methods across four well-studied protein families ($\beta$-lactamases, cytochrome c, GFP, and Ras). After determining optimal amino acid sequences and conducting side-chain homology modeling, we assess the functional relevance and evolutionary potential of the design, aiming to build confidence in deep generative approaches for protein design.
Our key contributions are: (1) We present a deep generative model pipeline for early-stage de novo protein design, with a robust evaluation protocol adaptable to a wider range of protein structure generation methods. (2) Despite sequence dissimilarity, generated samples show structural similarity to experimentally determined proteins in their families, conserving key residues linking to functions. (3) Structural phylogenetic analysis of both experimental and generated proteins reveals evolutionary relationships, showing the potential of structural phylogenetics. (4) Molecular dynamics (MD) simulations and protein-ligand docking show the dynamic stability and functional potential of generated samples. These samples form binding pockets akin to wild-type (WT) proteins, with conserved residues that bind family-specific ligands at low energies. For instance, generated KRas-like structures exhibit switch region flexibility and GTP/GDP-dependent conformational shifts similar to WT. (5) FM samples show greater flexibility, diversity, and novelty, while SM better captures conserved regions, offering insights for model selection and optimization in future work.
2 Background: Protein Backbone Generation
@@
In this section and Appendix A, we review the key concepts behind Yim et al. (2023) and Bose et al. (2023)’s approaches for backbone generation, using SM and FM on $\mathrm{SE}(3)$.
Protein backbone representations.
Molecules can be intuitively represented as atomic point clouds; however, for macromolecules, this approach often leads to high dimensionality and data sparsity. A more compact, fragment-based representation involves using backbone rigid groups from AlphaFold (Jumper et al., 2021). For each amino acid residue $i\in[1,N]$, its group consists of the backbone atoms $[\mathrm{N},\mathrm{C}_{\alpha},\mathrm{C},\mathrm{O}]_{i}$ (Figure 1A). The goal is to learn how the rigid transformation (or frame) $\bm{\mathsf{T}}_{i}$ acts on the idealized coordinates, so that the transformed coordinates match the actual coordinates. For simplicity, $\bm{\mathsf{T}}_{i}\in\mathrm{SE}(3)$ decomposes into a rotation matrix $\bm{\mathsf{R}}_{i}\in\mathrm{SO}(3)$ and a translation vector $\bm{\mathsf{x}}_{i}\in\mathbb{R}^{3}$. An extra dihedral angle $\psi_{i}\in\mathrm{SO}(2)$ is introduced between the $\mathrm{C}_{\alpha}$ and C bond to optimize the positioning of the backbone oxygen ($\mathrm{O}$).
$\mathrm{SE}(3)$ score matching.
Let $\mathbf{T}_{t}=\left[\bm{\mathsf{T}}_{1,t},\dots,\bm{\mathsf{T}}_{N,t}\right]$ be the manifold of $N$ frames at time $t$. The forward process gradually perturbs the distribution $\mathbf{T}_{0}\sim p_{0}$, pushing it toward a simple distribution $p_{T}$, such as a Gaussian, by introducing a stochastic process independently in rotation and translation spaces (Equation 4). The reverse process approximately reconstructs $p_{0}$ by iteratively calculating the gradient of the log probability density (or the score) $\nabla\log p$ along a reverse path (Equation 5 and Equation 6). Since this score function is typically intractable in practice, a network $s(\theta,t,\cdot)$ is trained to approximate the conditional score $\nabla\log p_{t|0}$ at time $t\sim\mathcal{U}(0,T)$ (Vincent, 2011).
$\mathrm{SE}(3)$ flow matching.
The flow from $p_{T}$ to $p_{0}$ is defined as a set of continuously differentiable, time-indexed vector fields $\{\mathbf{u}_{t}\}$, where FM learns the time evolution of $\mathbf{u}_{t}$ to gradually transport the data from $p_{T}$ to $p_{0}$. Because $\mathbf{u}_{t}$ is generally intractable, Lipman et al. (2022) introduced conditional FM, training a neural network $v(\theta,t,\cdot)$ to approximate $\mathbf{u}_{t|0}$ (Equation 13). When the optimal transport condition is satisfied, it provides a smooth and efficient transformation path, improving training stability and reducing computational complexity (Pooladian et al., 2023).
3 Evaluations of Generated Structures
Figure 1: (A) Protein backbone and dihedral angles $\psi$ and $\phi$. (B) Ramachandran plots showing $\psi$ and $\phi$ angle distributions for generated protein backbones vs. experimental proteins. The top-right inset shows conformationally favored (light) and disallowed (dark) angle regions. (C) Structural overlay of WT E. coli TEM1 (PDB: 1BTL; Jelsch et al. (1993); green) and its E166N acylated intermediate (PDB: 1FQG; Brown et al. (2009); white) with penicillin (PNM). (D) Structure of the WT E. caballus heart cytochrome c (PDB: 1HRC; (Dickerson et al., 1967)) with heme C (HEC). (E) Structure of A. victoria GFP and its chromophore (PDB: 4KW4; Barnard et al. (2014)). (F) Structure of GDP-bound H. sapiens KRas protein (PDB: 4OBE; (Hunter et al., 2014)).
@@ 3.1 Data
We generated monomeric proteins across a range of families, covering diverse fold types and functions. These proteins include both natural and engineered mutations while preserving conserved core functional regions.
$\beta$-lactamases.
$\beta$-lactamases are enzymes that deactivate $\beta$-lactam antibiotics by hydrolyzing their $\beta$-lactam ring, contributing significantly to bacterial resistance (Lee et al., 2016). Using $\beta$-lactamase inhibitors can restore antibiotic efficacy by blocking this reaction (Behzadi et al., 2020). Since their discovery, $\beta$-lactamase diversity has rapidly expanded due to the evolution of bacterial resistance, making them ideal for protein modeling studies. For this study, we gathered structural data on 1,578 unique monomeric $\beta$-lactamases and their variants across Ambler classes (A, B, C, and D) from the BLDB (Naas et al., 2017) and the Protein Data Bank (PDB) (Berman, 2000). Class A $\beta$-lactamases are the most prevalent, with conserved active-site residues Ser70, Glu166, and Asn170 coordinating the hydrolytic water for deacylation (Tooke et al., 2019; Brown et al., 2009). Figure 1C shows the wild-type $\beta$-lactamase TEM1 (1BTL) alongside its acylated E166N intermediate (1FQG).
Cytochrome c.
Figure 1D shows horse cytochrome c, where a hydrophobic shell surrounds the heme group, with only $\sim 7.5\%$ of the surface available for electron transfer (Bushnell et al., 1990). The hydrophobic environment and iron coordination by His18 and Met80 maintain a high redox potential ( 260 mV) (Salemme, 1977). Phosphorylation occurs at Thr28, Thr47, Tyr48, and Tyr97 (Hüttemann et al., 2011), while Lys72, Lys73, and Lys87 bind phospholipids (Kagan et al., 2009). The ATP-binding pocket includes Glu69, Asn70, Lys88, and Lys72, Lys86, Lys87 (McIntosh et al., 1996).
Green fluorescent proteins (GFP).
GFP, first isolated from Aequorea victoria, emits green fluorescence under specific wavelengths. Its core structure is an 11-strand $\beta$-barrel that encloses the chromophore (Figure 1E) (Remington, 2011). Various mutants have been engineered to enhance or modify its properties, including enhanced GFP (Cormack et al., 1996), superfolder GFP (Pédelacq et al., 2005), and color variants like YFP (Ormö et al., 1996) and BFP (Glaser et al., 2016). We collected structural data for 448 GFPs and variants from the PDB to explore this diversity.
Ras.
Ras proteins, a subgroup of the small GTPase superfamily, act as molecular switches, cycling between GTP-bound (active) and GDP-bound (inactive) states to regulate cell proliferation, differentiation, migration, and apoptosis (Ladygina et al., 2011; Simanshu et al., 2017; Weinmann & Ottow, 2007). They play a key role in signaling from the cell surface to downstream pathways. Mutations that keep Ras proteins in an active state drive excessive cell growth and malignancy, making Ras inhibition a promising cancer treatment strategy (Singh & Lingham, 2002). We focused on the most common cancer-related Ras proteins (HRas, KRas, and NRas) (Cox, 2002) and obtained 511 experimental structures from the PDB.
In human KRas (Figure 1F), the switch I and II regions form the key interface for interacting with effectors and regulators (Pantsar, 2020). These regions are highly flexible, with conformations depending on GTP or GDP binding. Cancer-related mutations frequently occur in the P-loop and switch II (Pantsar, 2020).
3.2 Training and Sampling
We trained the SM and FM (with OT) models on the described protein families, using pretrained weights from Yim et al. (2023) and Bose et al. (2023), with each model having $\sim 17$ million parameters. The sequence length distributions of the training proteins are shown in Figure 6. Each model generated 50 backbone structures with target lengths of 270 amino acids for $\beta$-lactamases, 113 for cytochrome c, 238 for GFP, and 170 for Ras proteins.
@@ 3.3 Backbone Dihedral Angles
Due to steric hindrance and spatial repulsion, not all backbone dihedral angles in proteins are physically feasible or energetically favorable. We analyzed the $\phi$ and $\psi$ distribution in the generated structures using Ramachandran plots (Ramachandran et al., 1963) (Figure 1B). Most data points fall within the allowed and favored regions, with only a few in the disallowed areas, and no significant geometric clashes or unreasonable conformations were observed.
The dihedral angle distributions of the generated structures align well with those of the experimental training proteins. For instance, cytochrome c structures have sparse points in the region $-180^{\circ}<\psi<-90^{\circ}$ and $45^{\circ}<\phi<180^{\circ}$, while GFP structures form four clusters there. SM samples are closer to the training data and concentrate in allowed regions but show less diversity than FM samples.
@@ 3.4 Conserved Residue Consistency
Figure 2: (A) Normalized evolutionary rates mapped onto structures, with black for rapidly evolving positions and orange for conserved ones. Experimental structures are highlighted with bold outlines, and highly conserved residues are marked in red (exposed) and green (buried). (B) Summary tree of two structural phylogenetic trees (Figure 7), constructed using $Q_{\text{score}}$ and the 3Di alphabet. Protein family members are grouped in semi-circular rings, with specific subtypes labeled in white. In $beta$-lactamases, different Ambler classes are shown with colored backgrounds. Colored branches and nodes indicate generated protein structures.
In protein families or across species, certain residues are highly conserved, typically to ensure structural stability and support proper folding and function.
Determining the optimal amino acid sequence.
Following Yim et al. (2023), we used ProteinMPNN (Dauparas et al., 2022) to predict ten sequences for each generated backbone. We then modeled these sequences with EMSfold (Rives et al., 2019) and compared them to the original backbone using the $\text{TM}_{\text{score}}$ (Zhang, 2005). The sequence with the highest $\text{TM}_{\text{score}}$ was chosen as the optimal match.
Identifying conserved residues.
The experimental protein sequences were aligned with each generated backbone’s optimal sequence using Clustal Omega (Sievers et al., 2011). Given these multiple sequence alignments (MSAs) and generated structures, ConSurf (Yariv et al., 2023) reconstructed phylogenetic trees and applied Rate4Site (Pupko et al., 2002) to calculate evolutionary rates for each position with an empirical Bayesian method (Ashkenazy et al., 2016; Mayrose, 2004). The optimal predicted sequences for the generated backbones showed conserved residues matching those in the experimental sequences (Figure 8). FM samples displayed greater diversity than SM samples, suggesting higher novelty, as noted in Section 3.5. Interestingly, this trend reverses in generated $\beta$-lactamases, possibly due to the greater structural variability in the experimental $\beta$-lactamase training set (Figure 2B), to which SM models are more sensitive. Section 3.6 further assesses the dynamic stability and ligand-binding properties of these generated samples.
In Figure 2A, similar to the experimental proteins, the generated structures have higher residue conservation near binding pockets. For instance, in cytochrome c (1HRC) and generated structures SM-1 and FM-0, conserved residues cluster around the central heme C binding site. A similar pattern is seen in 4OBE and the generated KRas proteins.
3.5 Structural Phylogenetics
In specific tasks, we aim to generate structures with dissimilar sequences while preserving similar functions. However, when sequence similarity falls below 30%, detecting homology and evolutionary relationships becomes challenging (Puente-Lelievre et al., 2023). Structural comparison, being more conserved (Illergård et al., 2009), is more effective for uncovering evolutionary connections (Flores et al., 1993; Puente-Lelievre et al., 2023; Moi et al., 2023).
Using $Q_{\text{score}}$ in structural phylogenetics.
Malik et al. (2020) proposed the $Q_{\text{score}}$ (Krissinel & Henrick, 2004) for structural phylogenetics, as it accommodates indels and combines alignment quality with length. $Q_{\text{score}}$ compares the positions of all $\mathrm{C}_{\alpha}$ atoms among $N_{\text{align}}$ comparable residues in pairwise comparisons. We construct structural phylogenetic trees using $1-Q_{\text{score}}$ as a distance measure, with higher values indicating greater structural similarity. For any two structures with $N_{1}$ and $N_{2}$ residues, we align and calculate $Q_{\text{score}}$ using TM-align (Zhang, 2005) as:
[…]
Using 3Di alphabet in structural phylogenetics.
van Kempen et al. (2023) developed Foldseek, which represents protein tertiary interactions as sequences over a 20-state 3D interaction (3Di) alphabet, simplifying structural alignments. This 3Di alphabet reduces false positives and increases information density by encoding conserved core regions more effectively. Leveraging Foldseek’s structural divergence metrics, Moi et al. (2023) constructed structural phylogenetic trees based on rigid body alignment, local alignment without superposition, and sequence alignment with structural alphabets, showing that these trees outperform traditional sequence-based trees across varied evolutionary timescales (Moi et al., 2023).

### v2

# Challenges and Guidelines in Deep Generative Protein Design: Four Case Studies

@@ Abstract
Deep generative models show promise for de novo protein design, yet reliably producing designs that are geometrically plausible, evolutionarily consistent, functionally relevant, and dynamically stable remains challenging. We present a deep generative modeling pipeline for early de novo design of monomeric proteins, based on Score Matching and Flow Matching. We apply this pipeline to four diverse protein families with an adaptable evaluation protocol. Generated structures display realistic, clash‐free conformations enriched with family-specific features, while the designed sequences preserve essential functional residues while retaining variability. Molecular dynamics and binding simulations show dynamic stability, with wild-type-like binding pockets that interact favorably with family-specific ligands. These results provide practical guidelines for integrating generative models into protein design workflows.
Generative Protein Design, Score Matching, Flow Matching, Structural Phylogenetics, Molecular Dynamics, Ligand Docking
@@ 1 Introduction
Protein functions and specificities are dictated by their complex structures. Over the past 60 years, we have progressed from viewing protein design as unattainable to achieving complete artificial design and synthesis (Korendovych & DeGrado, 2020; Huang et al., 2016), with expanding applications across industries (Arunachalam et al., 2021; Kingwell, 2024; Barclay & Acharya, 2023; Victorino da Silva Amatto et al., 2021; Ali et al., 2020). However, achieving atomic-level precision remains challenging due to the nonlinear complexity of folding and the sensitivity of the function to slight changes, still demanding significant resources.
@@
With the rapid growth of protein structure databases (Jamasb et al., 2024; Berman, 2000), various in silico de novo protein design methods, particularly deep generative models (Watson et al., 2023; Ingraham et al., 2023; Wu et al., 2022), have emerged. However, many methods generate structures that appear novel, diverse, and designable, yet few assess their functionality or evolutionary relevance, leaving the observations potentially as mere (Ji et al., 2023). Unclear biological functionality, unknown stability, and the lack of validation connecting these structures to known functions limit the broader application and advancement of these methods.
We tackle these challenges using a deep generative pipeline, based on diffusion-based Score Matching (SM) and Flow Matching (FM), for early de novo design of monomeric proteins. Guided by an adaptable evaluation protocol, we apply it to four protein families chosen for their diverse structural folds, functional roles, and rich annotations (Appendix C). Generated structures are realistic and clash-free, and structural phylogenetic analyses show that they capture family-specific features consistent with common ancestry and function. Designed sequences conserve key functional residues while allowing variability elsewhere, yielding low-similarity variants with similar functions. Molecular Dynamics (MD) simulations demonstrate dynamic stability of the designs, while forming wild-type (WT)-like binding pockets that favorably interact with family-specific ligands in docking studies. Together, these results offer practical guidelines for integrating generative models into protein design workflows.
2 Pipeline and Evaluation Protocol
2.1 Protein Backbone Generation
@@
In Appendix B, we review the key concepts behind the approaches of Yim et al. (2023) and Bose et al. (2023) for backbone generation, using SM and FM on $\mathrm{SE}(3)$.
2.2 Geometric Plausibility
Due to steric hindrance and spatial repulsion, not all backbone dihedral angles in proteins are physically feasible or energetically favorable. We analyzed the $\phi$ and $\psi$ distribution in the generated structures using Ramachandran plots (Ramachandran et al., 1963) (Figure 6).
2.3 Conserved Residue Consistency
In protein families or across species, certain residues are highly conserved, typically to ensure function but also structural stability and to support proper folding.
Determining the optimal amino acid sequence.
Following Yim et al. (2023), we used ProteinMPNN (Dauparas et al., 2022) to predict ten sequences for each generated backbone. We then modeled these sequences with EMSFold (Rives et al., 2019) and compared them to the generated backbone using the $\text{TM}_{\text{score}}$ (Zhang, 2005). The sequence with the highest $\text{TM}_{\text{score}}$ was chosen as the optimal match.
Identifying conserved residues.
Experimentally derived sequences were aligned with optimal sequences of generated backbones using Clustal Omega (Sievers et al., 2011). Given these alignments and generated structures, ConSurf (Yariv et al., 2023) reconstructed phylogenetic trees and applied Rate4Site (Pupko et al., 2002) to estimate per-residue evolutionary rates via an empirical Bayesian method (Ashkenazy et al., 2016; Mayrose, 2004).
2.4 Structural Phylogenetics
Certain applications require designing structures with low sequence similarity while retaining similar functions. However, when sequence identity falls below 30%, homology detection and evolutionary inference become challenging (Puente-Lelievre et al., 2023). Structural comparisons, which are more conserved, offer a more effective alternative (Illergård et al., 2009; Flores et al., 1993; Moi et al., 2023).
Using $Q_{\text{score}}$ in structural phylogenetics.
Malik et al. (2020) proposed the $Q_{\text{score}}$ (Krissinel & Henrick, 2004) for structural phylogenetics, which accommodates indels and combines alignment quality with length. It compares the positions of all $\mathrm{C}_{\alpha}$ atoms across $N_{\text{align}}$ comparable residues in pairwise comparisons. We construct structural phylogenetic trees using $1-Q_{\text{score}}$ as a distance measure, where higher values indicate greater structural similarity.
Using 3Di alphabet in structural phylogenetics.
van Kempen et al. (2023) developed Foldseek, which encodes protein tertiary interactions using a 20-state 3D interaction (3Di) alphabet to simplify structural alignments. This approach reduces false positives and increases information density by effectively encoding conserved core regions. Leveraging Foldseek’s divergence metrics, Moi et al. (2023) constructed structural phylogenetic trees based on rigid body alignment, local alignment without superposition, and sequence alignment with structural alphabets, showing that these trees outperform traditional sequence-based trees across varied evolutionary timescales (Moi et al., 2023).
2.5 Molecular Dynamics
Structures that appear reasonable may, in fact, be unstable due to molecular dynamics, water interactions, and entropy, which are not accounted for during generation. To assess the dynamic stability of these generated structures, we conducted MD simulations under physiological conditions and analyzed their time-dependent behavior.
Homology modeling of side-chains.
We used homology modeling to add side-chains to the protein backbone. After determining the optimal sequence (Section 2.3), we selected template proteins with at least 45% sequence identity and a $\text{TM}_{\text{score}}$ of 0.75 or higher using FoldSeek (van Kempen et al., 2023). Sequences were aligned with Clustal Omega, and MODELLER (Šali & Blundell, 1993) generated possible side-chain conformations using statistical potentials and rotamer libraries, with backbone fixed. We chose the final side-chain arrangement based on lowest energy and minimal steric clashes, verifying model quality with PROCHECK (Laskowski et al., 1993, 1996) and WHAT_CHECK (Hooft et al., 1996), discarding low-quality models. MODELLER’s modeling leverages (a) the input alignment to position side-chains, (b) template structures to set spatial restraints that mimic contacts, (c) knowledge-based rotamer distributions to favor typical angles, and (d) an optimization process.
Simulation.
Hydrogen atoms were added using Reduce2 from the computational crystallography toolbox (Grosse-Kunstleve et al., 2002), and MD simulations were performed with GROMACS (Abraham et al., 2015). Proteins were placed in an octahedral box with at least 15 Å̊$\mathring{\mathrm{A}}$ from the edges, and intermolecular interactions were modeled using the CHARMM36 force field (July 2022) (Vanommeslaeghe et al., 2009; Vanommeslaeghe & MacKerell, 2012; Yu et al., 2012). After vacuum energy minimization, the system was solvated, neutralized to 150 mM Na+ and Cl-, and minimized again. Then it was heated to 310 K under NVT [^1] Constant number of particles, volume, and temperature., equilibrated at 1 atm under NPT222constant number of particles, pressure, and temperature, and subjected to a 10 ns production run. Details are provided in Appendix I.
2.6 Protein-ligand Docking
AutoDock Vina (Trott & Olson, 2009; Eberhardt et al., 2021) was used to predict optimal binding modes between generated structures and their family-specific ligands. We performed blind docking, scanning the entire protein surface for potential ligand binding sites without prior knowledge of the pockets. The grid box covered the entire protein, and identical configurations were applied to both generated and experimentally derived structures to evaluate whether deviations in the generated samples fell within an acceptable range, thus assessing their functional viability. Details on simulation settings are provided in Section I.3.
3 Case Studies
@@ 3.1 Data
We assembled a dataset of monomeric proteins covering four families ($\beta$-lactamases, cytochrome c, GFP, and Ras), with varied fold types and functions, incorporating both natural and engineered mutations while retaining conserved core functional regions. Details are provided in Appendix C.
3.2 Backbone Generation
We trained the SM and FM (with Optimal Transport) models on these protein families, using pretrained weights from Yim et al. (2023) and Bose et al. (2023), with each model having $\sim 17$ million parameters. Each model generated 50 backbone structures, with target sequence lengths sampled from the distribution observed in the training data (Figure 5).
@@ 3.3 Backbone Dihedral Angles
In Figure 6, most data points fall within the allowed and favored regions, with few in the disallowed areas, and no significant geometric clashes or unreasonable conformations observed. The dihedral angle distributions of the generated structures align well with those of the experimentally derived proteins used for training. For instance, cytochrome c structures have sparse points in the region $-180^{\circ}<\psi<-90^{\circ}$ and $45^{\circ}<\phi<180^{\circ}$, while GFP structures form four clusters there. SM samples are closer to the training data and concentrate in allowed regions and show less diversity than FM samples.
@@ 3.4 Conserved Residue Consistency
The optimal sequences preserved conserved residues that were largely consistent with the experimentally derived data (Figure 8). In Figure 7, similar to the experimentally derived proteins, the generated structures have increased residue conservation around binding pockets (refer to Figure 4 and Figure 12). For instance, in cytochrome c (1HRC) and generated SM-1 and FM-0, conserved residues cluster near the central heme C binding site. A similar pattern appears in 4OBE and the generated KRas proteins.
3.5 Structural Phylogenetic Tree